# web_appendix

### v4

@@ 5.1 Notation and assumptions
Point 1 of Assumption 9 requires that $Y_{2}(D_{2}(z))-Y_{1}(D_{1}(z))$, units’ outcome evolutions in the counterfactual where their instrument does not change from period one to two, be mean independent of $Z_{2}$, conditional on $Z_{1}$ and $D_{1}$. Unlike Assumption 1, this condition imposes some restrictions on treatment effect heterogeneity, and the goal of conditioning on $D_{1}$ is to minimize the stringency of those restrictions. To see this, note that the two following conditions are sufficient for Point 1 of Assumption 9 to hold:
@@
(12) requires that $Y_{2}(D_{1}(z))-Y_{1}(D_{1}(z))$, units’ outcome evolutions in the counterfactual where their instrument and their treatment does not change from period one to two, be mean independent of $Z_{2}$, conditional on $Z_{1}$ and $D_{1}$. Thanks to the conditioning on $D_{1}$, (12) is a standard parallel trends assumption that does not impose any restriction on treatment effect heterogeneity, like Assumption 1. If $D_{1}$ was not conditioned upon, (12) would require parallel trends among units with different baseline treatments, which implicitly assumes homogeneous treatment effects over time, as discussed in Section 2. (13), on the other hand, is a restriction on treatment effect heterogeneity across units. Essentially, it requires that switching the treatment from $D_{1}(Z_{1})$ to $D_{2}(Z_{1})$, the natural treatment change happening over time even without any change in the instrument, has an effect on the outcome that is mean independent of $Z_{2}$ conditional on $Z_{1}$ and $D_{1}$. Importantly, note that Point 1 of Assumption 9 is placebo testable, by comparing the outcome evolutions of instrument-switchers and instrument-stayers before instrument-switchers experience a change of their instrument. Finally, Point 2 of Assumption 9 requires that units’ treatment evolutions under $Z_{1}$ be mean independent of $Z_{2}$, conditional on $Z_{1}$ and $D_{1}$. Because $D_{1}$ is conditioned upon, this parallel trends condition is equivalent to a sequential exogeneity assumption (see Robins, 1986; Bojinov et al., 2021).
Point 1 of Assumption 9 is related to identifying assumptions previously proposed in the literature. de Chaisemartin (2010) and Hudson et al. (2017) also consider IV-DID estimands, in classical designs with two periods and a binary instrument that turns on for some units at period two. Both papers introduce a “reduced-form” parallel trends assumption similar to Point 1 of Assumption 9, but without noting that it imposes restrictions on effects’ heterogeneity, even in the simple designs considered by those papers.
We also make the following assumptions.
Assumption 10
(Monotonicity and strictly positive first-stage) i) For all $(z,z^{\prime})\in\mathcal{Z}_{2}^{2}$, $z\geq z^{\prime}\Rightarrow D_{2}(z)\geq D_{2}(z^{\prime})$, and ii) $E(|D_{2}(Z_{2})-D_{2}(Z_{1})|)>0$.
i) is a monotonicity assumption similar to that in Imbens and Angrist (1994). It requires that increasing the period-two instrument weakly increases the period-two treatment. This condition is plausible when the instrument is taxes and the treatment is prices, as is the case in our application. ii) requires that the instrument has a strictly positive first stage.
Assumption 11
(Bounded instrument, Lipschitz and bounded reduced-form potential outcomes and potential treatments)
  1. 1.
    $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ are bounded subsets of $\mathbb{R}$.
  2. 2.
    For all $t\in\{1,2\}$ and for all $(z,z^{\prime})\in\mathcal{Z}_{t}^{2}$, there is a random variable $\overline{Y}\geq 0$ such that $|Y_{t}(D_{t}(z))-Y_{t}(D_{t}(z^{\prime}))|\leq\overline{Y}|z-z^{\prime}|$, with $\sup_{(z_{1},z_{2})\in\text{Supp}(Z_{1},Z_{2})}E[\overline{Y}|Z_{1}=z_{1},Z_{2}=z_{2}]<\infty$.
  3. 3.
    For all $t\in\{1,2\}$ and for all $(z,z^{\prime})\in\mathcal{Z}_{t}^{2}$, there is a random variable $\overline{D}\geq 0$ such that $|D_{t}(z)-D_{t}(z^{\prime})|\leq\overline{D}|z-z^{\prime}|$, with $\sup_{(z_{1},z_{2})\in\text{Supp}(Z_{1},Z_{2})}E[\overline{D}|Z_{1}=z_{1},Z_{2}=z_{2}]<\infty$.
Assumption 11 is an adaptation of Assumption 2 to the IV setting we consider in this section.
Assumption 12
(iid sample) We observe $(Y_{i,1},Y_{i,2},D_{i,1},D_{i,2},Z_{i,1},Z_{i,2})_{1\leq i\leq n}$, that are independent and identically distributed with the same probability distribution as $(Y_{1},Y_{2},D_{1},D_{2},Z_{1},Z_{2})$.
5.2 Target parameter
In this section, our target parameter is
$\delta_{IV}:=$ $E\left(\frac{|D_{2}(Z_{2})-D_{2}(Z_{1})|}{E(|D_{2}(Z_{2})-D_{2}(Z_{1})||SC=1)}\times\frac{Y_{2}(D_{2}(Z_{2}))-Y_{2}(D_{2}(Z_{1}))}{D_{2}(Z_{2})-D_{2}(Z_{1})}\middle|SC=1\right).$
$\delta_{IV}$ is a weighted average of the slopes of compliers-switchers’ period-two potential outcome functions, from their period-two treatment under their period-one instrument, to their period-two treatment under their period-two instrument. Slopes receive a weight proportional to the absolute value of compliers-switchers’ treatment response to the instrument change. $\delta_{IV}$ is just equal to the reduced-form WAS effect of the instrument on the outcome, divided by the first-stage WAS effect of the instrument on the treatment. With a binary instrument, such that $Z_{1}=0$ and $Z_{2}\in\{0,1\}$, our IV-WAS effect coincides with that identified in Corollary 2 of Angrist et al. (2000), in a cross-sectional IV model.
We could also consider a reduced-form AS divided by a first-stage AS. The resulting target is a weighted average of the slopes $\frac{Y_{2}(D_{2}(Z_{2}))-Y_{2}(D_{2}(Z_{1}))}{D_{2}(Z_{2})-D_{2}(Z_{1})}$, with weights proportional to $\frac{D_{2}(Z_{2})-D_{2}(Z_{1})}{Z_{2}-Z_{1}}$. It seems more natural to us to weight compliers-switchers’ slopes by the absolute value of their first-stage than by the slope of their first-stage. [^6] If the first-stage effect is homogenous and linear, the weights in the IV-AS effect reduce to one, and one recovers a standard AS effect. However, linearity and homogeneity of the first-stage effect are strong assumptions.
5.3 Identification
Let $S^{I}=1\{Z_{2}-Z_{1}\neq 0\}$, $S^{I}_{+}=1\{Z_{2}-Z_{1}>0\}$, and $S^{I}_{-}=1\{Z_{2}-Z_{1}<0\}$.
Assumption 13
(Support conditions for IV-WAS identification)
  1. 1.
    $0<P(S^{I}_{+}=1)$, and $0<P(S^{I}_{+}=1|Z_{1},D_{1})$ implies that $0<P(S^{I}=0|Z_{1},D_{1})$.
  2. 2.
    $0<P(S^{I}_{-}=1)$, and $0<P(S^{I}_{-}=1|Z_{1},D_{1})$ implies that $0<P(S^{I}=0|Z_{1},D_{1})$.
Theorem 6
If Assumptions 9-11 and 13 hold,
$\delta_{IV}=$ $\frac{E\left[\text{sgn}(\Delta Z)\left(\Delta Y-E(\Delta Y|Z_{1},D_{1},S^{I}=0)\right)\right]}{E\left[\text{sgn}(\Delta Z)\left(\Delta D-E(\Delta D|Z_{1},D_{1},S^{I}=0)\right)\right]}$ (14)
$=$ $\frac{E\left[\text{sgn}(\Delta Z)\Delta Y\right]-E\left[\Delta Y\frac{P(S^{I}_{+}=1|Z_{1},D_{1})-P(S^{I}_{-}=1|Z_{1},D_{1})}{P(S^{I}=0|Z_{1},D_{1})}P(S^{I}=0)\middle|S^{I}=0\right]}{E\left[\text{sgn}(\Delta Z)\Delta D\right]-E\left[\Delta D\frac{P(S^{I}_{+}=1|Z_{1},D_{1})-P(S^{I}_{-}=1|Z_{1},D_{1})}{P(S^{I}=0|Z_{1},D_{1})}P(S^{I}=0)\middle|S^{I}=0\right]}.$ (15)
@@ 5.3 Identification
The regression-based (resp. propensity-score-based) estimand identifying $\delta_{IV}$ is just equal to the regression-based (resp. propensity-score-based) estimand identifying the reduced-form WAS effect of the instrument on the outcome controlling for $D_{1}$, divided by the regression-based (resp. propensity-score-based) estimand identifying the first-stage WAS effect controlling for $D_{1}$.
5.4 Estimation and inference
Let
$\widehat{\delta}^{r}_{IV}=$ $\frac{\frac{1}{n}\sum_{i=1}^{n}\text{sgn}(\Delta Z_{i})\left(\Delta Y_{i}-\widehat{E}(\Delta Y|Z_{1}=Z_{i,1},D_{1}=D_{i,1},S^{I}=0)\right)}{\frac{1}{n}\sum_{i=1}^{n}\text{sgn}(\Delta Z_{i})\left(\Delta D_{i}-\widehat{E}(\Delta D|Z_{1}=Z_{i,1},D_{1}=D_{i,1},S^{I}=0)\right)},$ (16)
where $\widehat{E}(\Delta Y|Z_{1},D_{1},S^{I}=0)$ and $\widehat{E}(\Delta D|Z_{1},D_{1},S^{I}=0)$ are series estimators of $E(\Delta Y|Z_{1},D_{1},S^{I}=0)$ and $E(\Delta D|Z_{1},D_{1},S^{I}=0)$ defined analogously to the series estimator in Section 3.3.
Let us define
$\widehat{\delta}^{ps}_{IV}=$ $\frac{\frac{1}{n}\sum_{i=1}^{n}\text{sgn}(\Delta Z_{i})\Delta Y_{i}-\frac{1}{n}\sum_{i:S^{I}_{i}=0}\Delta Y_{i}\frac{\widehat{P}(S^{I}_{+}=1|Z_{1}=Z_{i1},D_{1}=D_{i1})-\widehat{P}(S^{I}_{-}=1|Z_{1}=Z_{i1},D_{1}=D_{i1})}{\widehat{P}(S^{I}=0|Z_{1}=Z_{i1},D_{1}=D_{i1})}}{\frac{1}{n}\sum_{i=1}^{n}\text{sgn}(\Delta Z_{i})\Delta D_{i}-\frac{1}{n}\sum_{i:S^{I}_{i}=0}\Delta D_{i}\frac{\widehat{P}(S^{I}_{+}=1|Z_{1}=Z_{i1},D_{1}=D_{i1})-\widehat{P}(S^{I}_{-}=1|Z_{1}=Z_{i1},D_{1}=D_{i1})}{\widehat{P}(S^{I}=0|Z_{1}=Z_{i1},D_{1}=D_{i1})}},$ (17)
where $\widehat{P}(S^{I}_{+}=1|Z_{1},D_{1})$ (resp. $\widehat{P}(S^{I}_{-}=1|Z_{1},D_{1})$, $\widehat{P}(S^{I}=0|Z_{1},D_{1})$) is a series logistic regression estimator of $P(S^{I}_{+}=1|Z_{1},D_{1})$ (resp. $P(S^{I}_{-}=1|Z_{1},D_{1})$, $P(S^{I}=0|Z_{1},D_{1})$) defined analogously to the series logistic regression estimators in Section 4.3.
For any variable $X$, let
$\delta_{X}$ $=E\left[\text{sgn}(\Delta Z)\left(\Delta X-E(\Delta X|Z_{1},D_{1},S^{I}=0)\right)\right]$
$\psi_{X}$ $=\frac{1}{E(\left|\Delta Z\right|)}\left\{\left(S^{I}_{+}-S^{I}_{-}-E(S^{I}_{+}-S^{I}_{-}|Z_{1},D_{1})\frac{(1-S^{I})}{E(1-S^{I}|Z_{1},D_{1})}\right)\right.$
×$\qquad\times(\Delta X-E(\Delta X|Z_{1},D_{1},S^{I}=0))-\delta_{X}\left|\Delta Z\right|\bigg{\}}.$
Then, let
$\psi_{IV}$ $=\frac{\psi_{Y}-\delta_{IV}\psi_{D}}{\delta_{D}}.$
Under technical conditions similar to those in Assumptions 6 and 8, one can show that
$\sqrt{n}\left(\widehat{\delta}^{r}_{IV}-\delta_{IV}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{IV})),$
$\sqrt{n}\left(\widehat{\delta}^{ps}_{IV}-\delta_{IV}\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,V(\psi_{IV})).$
6 Extensions
@@
In this section, we return to the case where the treatment, rather than an instrument, satisfies a parallel-trends condition. Combining the extensions below with the IV case is possible.
6.1 More than two time periods
In this section, we assume the representative unit is observed at $T>2$ time periods. Let $(D_{1},...,D_{T})$ denote the unit’s treatments and $\mathcal{D}_{t}=\text{Supp}(D_{t})$ for all $t\in\{1,...,T\}$. For any $t\in\{1,...,T\}$, and for any $d\in\mathcal{D}_{t}$ let $Y_{t}(d)$ denote the unit’s potential outcome at period $t$ with treatment $d$. Finally, let $Y_{t}$ denote their observed outcome at $t$. For any $t\in\{2,...,T\}$, let $S_{t}=1\{D_{t}\neq D_{t-1}\}$ be an indicator equal to 1 if the unit’s treatment switches from period $t-1$ to $t$. Let also $S_{+,t}=1\{D_{t}>D_{t-1}\}$ and $S_{-,t}=1\{D_{t}<D_{t-1}\}$. We assume that the assumptions made in the paper, rather than just holding for $t=1$ and $t=2$, actually hold for all pairs of consecutive time periods $(t-1,t)$. For instance, we replace Assumption 1 by:
Assumption 14
(Parallel trends) For all $t\geq 2$, for all $d\in\mathcal{D}_{t-1}$, $E(\Delta Y_{t}(d)|D_{t-1}=d,D_{t})=E(\Delta Y_{t}(d)|D_{t-1}=d)$.
Assumption 14 requires that $E(\Delta Y_{t}(d)|D_{t-1}=d,D_{t}=d^{\prime})$ be constant across $d^{\prime}$: groups of units with the same period-$t-1$ treatment but different period-$t$ treatments all have the same expected outcome evolution in the counterfactual where their period-$t-1$ treatment would not have changed. Importantly, note that because Assumption 14 is conditional on $D_{t-1}$, it cannot be “chained” across pairs of time periods: it requires parallel trends over pairs of consecutive time periods, not over the entire duration of the panel. To preserve space, we do not restate our other assumptions with more than two periods.
Let
$\delta_{1,t}$ $=E\left(\frac{Y_{t}(D_{t})-Y_{t}(D_{t-1})}{D_{t}-D_{t-1}}\middle|S_{t}=1\right),$
$\delta_{2,t}$ $=\frac{E\left(\text{sgn}(D_{t}-D_{t-1})(Y_{t}(D_{t})-Y_{t}(D_{t-1}))\right)}{E(|D_{t}-D_{t-1}|)}.$
Let
$\delta^{T\geq 3}_{1}$ $=\sum_{t=2}^{T}\frac{P(S_{t}=1)}{\sum_{k=2}^{T}P(S_{k}=1)}\delta_{1,t},$
$\delta^{T\geq 3}_{2}$ $=\sum_{t=2}^{T}\frac{E(|\Delta D_{t}|)}{\sum_{k=2}^{T}E(|\Delta D_{k}|)}\delta_{2,t}$
be generalizations of the AS and WAS effects to applications with more than two periods. Note that in line with the spirit of the two effects, we propose different weights to aggregate the AS and WAS across time periods. For the AS, the weights are just proportional to the proportion of switchers between $t-1$ and $t$. For the WAS, the weights are proportional to the average absolute value of the treatment switch from $t-1$ to $t$.
Theorem 7
If Assumption 14 and generalizations of Assumptions 2-5 to more than two periods hold,
$\delta^{T\geq 3}_{1}=$ $\sum_{t=2}^{T}\frac{P(S_{t}=1)}{\sum_{k=2}^{T}P(S_{k}=1)}E\left(\frac{\Delta Y_{t}-E(\Delta Y_{t}|D_{t-1},S_{t}=0)}{\Delta D_{t}}\middle|S_{t}=1\right).$
Theorem 8
If Assumption 14 and generalizations of Assumptions 2 and 7 to more than two periods hold,
$\delta^{T\geq 3}_{2}=$ $\sum_{t=2}^{T}\frac{E(|\Delta D_{t}|)}{\sum_{k=2}^{T}E(|\Delta D_{k}|)}\frac{E\left(\text{sgn}(\Delta D_{t})\left(\Delta Y_{t}-E(\Delta Y_{t}|D_{t-1},S_{t}=0)\right)\right)}{E(|\Delta D_{t}|)}$
$=$ $\sum_{t=2}^{T}\frac{E(|\Delta D_{t}|)}{\sum_{k=2}^{T}E(|\Delta D_{k}|)}\frac{E\left[\text{sgn}(\Delta D_{t})\Delta Y_{t}\right]-E\left[\Delta Y_{t}\frac{P(S_{+,t}=1|D_{t-1})-P(S_{-,t}=1|D_{t-1})}{P(S_{t}=0|D_{t-1})}P(S_{t}=0)\middle|S_{t}=0\right]}{E(|\Delta D_{t}|)}.$
Theorems 7 and 8 are straightforward generalizations of Theorems 1 and 4 to settings with more than two time periods.
Let
$\psi_{1,t}=$ $\frac{1}{E(S_{t})}\left\{\left(\frac{S_{t}}{\Delta D_{t}}-E\left(\frac{S_{t}}{\Delta D_{t}}\middle|D_{t-1}\right)\frac{(1-S_{t})}{E[1-S_{t}|D_{t-1}]}\right)[\Delta Y_{t}-E(\Delta Y_{t}|D_{t-1},S_{t}=0)]-\delta_{1,t}S_{t}\right\},$
$\psi_{2,t}=$ $\frac{1}{E(|\Delta D_{t}|)}\left\{\left(S_{+,t}-S_{-,t}-E(S_{+,t}-S_{-,t}|D_{t-1})\frac{(1-S_{t})}{E(1-S_{t}|D_{t-1})}\right)(\Delta Y_{t}-E(\Delta Y_{t}|D_{t-1},S_{t}=0))-\delta_{2,t}|\Delta D_{t}|\right\}.$
After some algebra, one can show that the influence function of the AS estimator with several periods is
$\psi^{T\geq 3}_{1}$ $:=\frac{\sum_{t=2}^{T}(P(S_{t}=1)\psi_{1,t}+(\delta_{1,t}-\delta^{T\geq 3}_{1})(S_{t}-P(S_{t}=1)))}{\sum_{t=2}^{T}P(S_{t}=1)},$ (18)
while the influence function of the WAS estimators with several periods is
$\psi^{T\geq 3}_{2}$ $:=\frac{\sum_{t=2}^{T}E(\left|\Delta D_{t}\right|)\psi_{2,t}+(\delta_{2,t}-\delta^{T\geq 3}_{2})(\left|\Delta D_{t}\right|-E(\left|\Delta D_{t}\right|))}{\sum_{t=2}^{T}E(|\Delta D_{t}|)}.$ (19)
Importantly, those influence functions allow the unit’s treatments and outcomes to be arbitrarily serially correlated.
6.2 Placebo tests
With several time periods, one can test the following condition, which is closely related to Assumption 14:
Assumption 15
(Testable parallel trends) For all $t\geq 3,t\leq T$, for all $d\in\mathcal{D}_{t-1}$, $E(\Delta Y_{t-1}(d)|D_{t-2}=D_{t-1}=d,D_{t})=E(\Delta Y_{t-1}(d)|D_{t-2}=D_{t-1}=d)$.
To test that condition, one can compute a placebo version of the estimators described in the previous subsection, replacing $\Delta Y_{t}$ by $\Delta Y_{t-1}$, and restricting the sample, for each pair of consecutive time periods $(t-1,t)$, to units whose treatment did not change between $t-2$ and $t-1$. Thus, the placebo compares the average $\Delta Y_{t-1}$ of the $t-1$-to-$t$ switchers and stayers, restricting attention to $t-2$-to-$t-1$ stayers. If one finds that from $t-2$-to-$t-1$, $t-1$-to-$t$ switchers and stayers are on parallel trends, this lends credibility to Assumption 14.
Assumption 14 can only be placebo tested among $t-2$-to-$t-1$ stayers. Then, as a robustness check one may restrict the estimation of $\delta_{1}$ and $\delta_{2}$ to $t-2$-to-$t-1$ stayers, to ensure that effects are only estimated in a subsample for which the identifying assumption can be placebo tested. The resulting estimator relies on the following identifying assumption:
$\forall t\geq 3,t\leq T,d\in\mathcal{D}_{t-1}:\leavevmode\nobreak\ E(\Delta Y_{t}(d)|D_{t-2}=D_{t-1}=d,D_{t})=E(\Delta Y_{t}(d)|D_{t-2}=D_{t-1}=d),$
the exact analogue of Assumption 15 but one period ahead.
6.3 Estimators robust to dynamic effects up to a pre-specified treatment lag.
Importantly, the robustness check in the previous section also yields an estimator robust to dynamic effects up to one treatment lag. If units’ current and first treatment lag affect their current outcome, our $t-1$-to-$t$ estimators in the subsample of $t-2$-to-$t-1$ stayers are unbiased for effects of the current treatment on the outcome under the following assumption:
$\forall t\geq 3,t\leq T,d\in\mathcal{D}_{t-1}:\leavevmode\nobreak\ E(Y_{t}(d,d)-Y_{t-1}(d,d)|D_{t-2}=D_{t-1}=d,D_{t})=E(Y_{t}(d,d)-Y_{t-1}(d,d)|D_{t-2}=D_{t-1}=d).$
Similarly, if one wants to allow for effects of the first and second treatment lags on the outcome, one just needs to restrict the estimation sample to $t-3$-to-$t-1$ stayers. However, the more robustness to dynamic effects one would like to have, the smaller the estimation sample becomes.

### v5

@@ 5.1 Instrumental-variable estimators
Point 1 of Assumption 9 requires that $Y_{2}(D_{2}(z))-Y_{1}(D_{1}(z))$, units’ outcome evolutions in the counterfactual where their instrument does not change from period one to two, be mean independent of $Z_{2}$, conditional on $Z_{1}$ and $D_{1}$. This condition imposes some restrictions on treatment effect heterogeneity, and the goal of conditioning on $D_{1}$ is to alleviate those restrictions. To see this, note that the two following conditions are sufficient for Point 1 of Assumption 9 to hold:
@@
(2) requires that $Y_{2}(D_{1}(z))-Y_{1}(D_{1}(z))$, units’ outcome evolutions in the counterfactual where their instrument and their treatment does not change from period one to two, be mean independent of $Z_{2}$, conditional on $Z_{1}$ and $D_{1}$. Thanks to the conditioning on $D_{1}$, (2) is a standard parallel-trends assumption that does not impose any restriction on treatment effect heterogeneity, like Assumption 1. If $D_{1}$ was not conditioned upon, (2) would require parallel trends among units with different baseline treatments, which implicitly assumes homogeneous treatment effects over time, as discussed in Section 2.4. (3), on the other hand, is a restriction on treatment effect heterogeneity across units. It requires that switching the treatment from $D_{1}(Z_{1})$ to $D_{2}(Z_{1})$, the natural treatment change happening over time even without any change in the instrument, has an effect on the outcome that is mean independent of $Z_{2}$ conditional on $Z_{1}$ and $D_{1}$. de Chaisemartin (2010) and Hudson et al. (2017) also consider IV-DID estimands, in classical designs with two periods and a binary instrument that turns on for some units at period two. Both papers introduce a “reduced-form” parallel-trends assumption similar to Point 1 of Assumption 9, but without noting that it imposes restrictions on effects’ heterogeneity, even in the simple designs considered by those papers. [^7] The realization that the “reduced-form” parallel-trends assumption in de Chaisemartin (2010) and Hudson et al. (2017) actually restricts effect heterogeneity explains the apparent discrepancy between those two papers and Blundell and Dias (2009) and de Chaisemartin and D’Haultfœuille (2018), two papers that show that IV-DIDs restrict treatment-effect heterogeneity under the standard parallel-trends assumption on the untreated outcome. Turning to Point 2 of Assumption 9, this condition requires that units’ treatment evolutions under $Z_{1}$ be mean independent of $Z_{2}$, conditional on $Z_{1}$ and $D_{1}$. Because $D_{1}$ is conditioned upon, this parallel trends condition is equivalent to a sequential exogeneity assumption (see Robins, 1986; Bojinov et al., 2021).
Let us consider the parameter
$\delta_{\mathsf{IV}}:=$ $E\left(\frac{|D_{2}(Z_{2})-D_{2}(Z_{1})|}{E(|D_{2}(Z_{2})-D_{2}(Z_{1})||SC=1)}\times\frac{Y_{2}(D_{2}(Z_{2}))-Y_{2}(D_{2}(Z_{1}))}{D_{2}(Z_{2})-D_{2}(Z_{1})}\middle|SC=1\right).$
$\delta_{\mathsf{IV}}$ is a weighted average of the slopes of compliers-switchers’ period-two potential outcome functions, from their period-two treatment under their period-one instrument, to their period- two treatment under their period-two instrument. We show in Section LABEL:appendixsec:IV of the web appendix that under Assumption 9 and a monotonicity condition as in Imbens and Angrist (1994),
$\delta_{\mathsf{IV}}=\frac{E\left[\left(S^{I}_{+}-S^{I}_{-}-\gamma^{I}_{0}(S^{I},Z_{1},D_{1})\right)(\Delta Y-\mu^{Y}_{0}(Z_{1},D_{1}))\right]}{E\left[\left(S^{I}_{+}-S^{I}_{-}-\gamma^{I}_{0}(S^{I},Z_{1},D_{1})\right)(\Delta D-\mu^{D}_{0}(Z_{1},D_{1}))\right]},$
where $S^{I}=1\{Z_{2}-Z_{1}\neq 0\}$, $S^{I}_{+}=1\{Z_{2}-Z_{1}>0\}$, $S^{I}_{-}=1\{Z_{2}-Z_{1}<0\}$ and
$\mu^{U}_{0}(z,d):=E[\Delta U|S^{I}=0,Z_{1}=z,D_{1}=d],\quad U\in\{D,Y\},$
$\gamma^{I}_{0}(s,z,d):=\frac{E(S^{I}_{+}|Z_{1}=z,D_{1}=d)-E(S^{I}_{-}|Z_{1}=z,D_{1}=d)}{E(1-S^{I}|Z_{1}=z,D_{1}=d)}(1-s).$
@@
The estimand identifying $\delta_{\mathsf{IV}}$ is equal to the estimand identifying the reduced-form WAS effect of the instrument on the outcome controlling for $D_{1}$, divided by the estimand identifying the first-stage WAS effect of the instrument on the treatment controlling for $D_{1}$.
5.2 Estimators with more than two time periods
@@
In this section and the next, we return to the case where the treatment, rather than an instrument, satisfies a parallel-trends condition. This section sketches the extension of our estimators to datasets with three periods or more. For further details, see Section LABEL:appendixsec:Tgeq3 of the web appendix.
With $T>2$ time periods, we propose a simple extension of our estimators. Let $\widehat{\delta}_{1,t,\mathsf{DR}}$ and $\widehat{\delta}_{2,t,\mathsf{DR}}$ be defined exactly as $\hat{\delta}_{1,\mathsf{DR}}$ and $\hat{\delta}_{2,\mathsf{DR}}$ but on a pair of consecutive time periods $(t-1,t)$. Then, we consider a weighted average across $t$ of $\widehat{\delta}_{1,t,\mathsf{DR}}$, with weights proportional to the proportion of switchers from $t-1$ to $t$, thus yielding an AS estimator with more than two periods, denoted $\widehat{\delta}^{T\geq 3}_{1}$. We also consider a weighted average across $t$ of $\widehat{\delta}_{2,t,\mathsf{DR}}$, with weights proportional to the average absolute treatment change from $t-1$ to $t$, thus yielding a WAS estimator with more than two periods, denoted $\widehat{\delta}^{T\geq 3}_{2}$.
Those estimators are consistent under the following parallel-trends assumption: $t-1$-to-$t$ switchers and stayers with the same $t-1$ treatment would have experienced the same average $t-1$-to-$t$ outcome evolutions if switchers had not switched. We propose “pre-trends” or “placebo” estimators, to assess the plausibility of that assumption. The placebo AS and WAS estimators are analogous to $\widehat{\delta}^{T\geq 3}_{1}$ and $\widehat{\delta}^{T\geq 3}_{2}$, but they compare the $t-2$-to-$t-1$ outcome evolutions of $t-1$-to-$t$ switchers and stayers, restricting the sample to $t-2$-to-$t-1$ stayers to avoid that those comparisons be contaminated by the treatment’s effect. If one finds that from $t-2$-to-$t-1$, $t-1$-to-$t$ switchers and stayers are on parallel trends, this lends credibility to the parallel-trends assumption underlying $\widehat{\delta}^{T\geq 3}_{1}$ and $\widehat{\delta}^{T\geq 3}_{2}$.
With $T>2$ time periods, the assumption that units’ potential outcomes does not depend on their past treatments is of essence. If there are dynamic treatment effects, $\widehat{\delta}^{T\geq 3}_{1}$ and $\widehat{\delta}^{T\geq 3}_{2}$ may be biased. Those estimators compare the outcome evolutions of $t-1$-to-$t$ switchers and stayers. But there may be, say, $t-1$-to-$t$ stayers whose treatment changed from $t-2$ to $t-1$. If lagged treatments affect units’ current outcome, that change could still affect the $t-1$-to-$t$ outcome evolution of those stayers, thus leading them to violate the parallel-trends assumption underlying $\widehat{\delta}^{T\geq 3}_{1}$ and $\widehat{\delta}^{T\geq 3}_{2}$. To mitigate this concern, we propose the following robustness check. One can recompute $\widehat{\delta}^{T\geq 3}_{1}$ and $\widehat{\delta}^{T\geq 3}_{2}$, restricting, for each pair of consecutive periods $(t-1,t)$, the estimation sample to $t-2$-to-$t-1$ stayers, as in the placebo analysis. We show that the resulting estimator is robust to dynamic effects up to one treatment lag. Similarly, if one wants to allow for effects of the first and second treatment lags on the outcome, one just needs to restrict the estimation sample to $t-3$-to-$t-1$ stayers. However, the more robustness to dynamic effects one would like to have, the smaller the estimation sample becomes.
5.3 Identification and estimation with covariates
Our methods can be generalized to incorporate covariates. Consider for simplicity the case with two periods, $T=2$, where $X_{t}$ is a $c$-dimensional vector of covariates. We show in Section LABEL:appendixsec:covariates of the web appendix that the AS and WAS can be identified and estimated under an alternative parallel trends assumption that holds conditional on the first-period covariates $X_{1}$ (Assumption LABEL:hyp:parallel_trendsX). Because our estimation and inference methods are fully nonparametric, they can be readily applied to the case with covariates after redefining the nuisance functions to depend on the covariates, for instance, $\mu_{0}(x,d)=E(\Delta Y|X_{1}=x,D_{1}=d,S=0)$ and so on. The main difference is that with covariates, the nuisance functions are nonparametric objects of dimension $c+1$, so classical nonparametric estimators, such as series estimators, may not satisfy the convergence rate requirements for asymptotic normality of the AS and WAS estimators. Then, it may be preferable to use machine-learning estimators to estimate those nuisance functions.